A man is equally likely to choose any one of the three routes A, B, C from his office,

and his choice of route is not influenced by the weather. If the weather is dry the probabilities

of his being late by routes A, B, C are respectively 1/20, 1/10, 1/8. He sets out on a dry day and reaches his office late. What is the probability that the route chosen was C?


1
Expert's answer
2020-12-15T14:13:25-0500

This is a conditional probability case. let P(L) denote the probability of being late and P(C) be the probability of using route C.

"P(C|L)=\frac{P(L\bigcap C)}{P(L)}"

"P(L)=\frac{1}{3}*(\frac{1}{20}+\frac{1}{10}+\frac{1}{8})"

="\frac {11}{120}"

"P(L\bigcap C)=\frac{1}{3}*\frac{1}{8}" ="\frac{1}{24}"

"P(C|L)=\frac{1}{24} \div\frac{11}{120}"

="\frac{5}{11}"